OpenStudy (anonymous):

Which expression defines the given series for seven terms? 15 + 19 + 23 + . . . Using the [E]

OpenStudy (solomonzelman):

do you mean using \(\large\color{blueviolet}{ \displaystyle \sum_{ }^{ } }\) ?

OpenStudy (solomonzelman):

your difference is 4 (that is what you are adding every time to find the next term) your first term term is 15. \(\LARGE\color{blueviolet}{ \displaystyle \sum_{{\large n{\Large =}{\rm start~plugging~from } }}^{ {\large {\rm number~of~terms} }} ~ {\rm the~pattern}}\)

OpenStudy (solomonzelman):

what can you tell me so far ?

OpenStudy (anonymous):

15+19+23+ ... ( 27,31,35)

OpenStudy (solomonzelman):

you don;t need to find those terms

OpenStudy (solomonzelman):

I will give you an example of how to write a pattern. Sequence: \(\large\color{blueviolet}{ \displaystyle 4,~~~9,~~~14,~~~19,~~~...}\) write the expression that will represent 100 terms. You know that \(\large\color{blueviolet}{ \displaystyle a_1=4}\) \(\large\color{blueviolet}{ \displaystyle d=5}\) \(\large\color{blueviolet}{ \displaystyle \sum_{ n=0 }^{ 99 } ~(5n+4)}\)

OpenStudy (solomonzelman):

you plug in n=0 for the first term (and get 4) plug in n=1 for the second term (and get 9) and on... you will have n=99 for 100th term. so it grows by 5, or in other words, the slope is 5. And the a(1)=4, in other words y-intercept is 4.

OpenStudy (solomonzelman):

is this example making sense ?

OpenStudy (anonymous):

so d is the number of how many times is adding up ?